# The quantum metric of electrons with spin-momentum locking

### v2

@@
We have finally performed a similar scaling analysis of the anomalous planar Hall resistance to account for the presence of disorder-mediated effects beside the intrinsic Berry curvature-induced contribution. The linear dependence of the anomalous planar Hall conductivity on the mobility $\mu_{2}$ (Fig. 4A) demonstrates the existence of a skew scattering contribution $\sigma_{\textrm{sk}}$ that is absent in a rotational symmetric Rashba two-dimensional electron system  (?) but is activated by the nonvanishing Berry curvature triple due to hexagonal warping effects  (?). In addition, the $\tau^{0}$ term due to the combined action of side-jump and intrinsic Berry curvature varies nonmonotonically as the magnetic field is increased. This behavior is correctly captured by a theory calculation of the Berry curvature-mediated anomalous planar Hall response (Fig. 4D), which thus proves that the intrinsic contribution dominates the $\tau$-independent anomalous planar Hall conductivity. Therefore, the intrinsic contributions to the anomalous planar Hall effect and nonlinear in-plane magnetoresistance provide us with direct access to the Berry curvature and the quantum metric of the 111-oriented LaAlO3/SrTiO3 interface and allow us to probe electronic dynamics associated with its full quantum geometric tensor. This unique property marks the difference of 111-oriented LaAlO3/SrTiO3 with respect to $\mathcal{P}\mathcal{T}$-symmetric materials like topological antiferromagnets  (?, ?), where the quantum geometric tensor is purely real and accessible only via the nonlinear response.
Outlook
@@
The analysis of the conductivity dependence on the scattering time allowed us to identify the intrinsic contribution of the quantum metric and Berry curvature to the electronic magnetotransport. While the Berry curvature is enabled by the trigonal symmetry of the 111-oriented interface, the quantum metric is rotationally invariant and is not contingent on specific symmetry requirements beside the broken inversion symmetry. Our findings, therefore, do not only demonstrate the existence of a quantum metric associated with the spin-momentum locking but unlock new directions for investigating and functionalizing the quantum geometry in a vast class of nonmagnetic materials. The QMMR is not limited to oxide heterostructures  (?, ?) but is expected for all materials with spin-locked electronic bands such as semiconductors  (?, ?, ?, ?), surface states of heavy metals  (?), magnetic and nonmagnetic interfaces  (?, ?, ?), and even policrystalline samples. Besides magnetotransport, the quantum metric of the spin-momentum locking is, in principle, accessible to THz photocurrents and spectroscopic tools in time-reversal symmetric conditions   (?, ?). In a technological perspective, integrating the spin-momentum locking with magnetic and ferroelectric thin films could lead to the on-demand, reversible, and nonvolatile control of the physical effects emerging from the quantum metric. Moreover, combining spintronics with the quantum metric may reveal unforeseen disorder-free mechanisms to interconvert charge and spin-orbital currents  (?, ?, ?). This study discloses therefore a whole range of possibilities to harness quantum metric effects in condensed matter and ultimately inspire technologies based on the geometry of the electronic wavefunctions.
@@ Acknowledgments
We thank Jean-Marc Triscone, Marc Gabay, Ajit Srivastava, Edouard Lesne, and Graham Kimbell for fruitful discussions. We thank Marco Lopes and Jean-Luc Lorenzoni for technical support. We thank Alberto Morpurgo and Ignacio Gutiérrez for contributing to the sample fabrication. Funding This work was supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract no. MB22.00071, by the Gordon and Betty Moore Foundation (grant no. 332 GBMF10451 to A.D.C.), by the European Research Council (ERC), and by the Dutch Research Council (NWO) as part of the VIDI (project 016.Vidi.189.061 to A.D.C.). M.T.M and M.C. acknowledge support from the EU Horizon 2020 research and innovation program under Grant Agreement No. 964398 (SUPERGATE). M.T.M. and C.O. acknowledge support from the MAECI project ”ULTRAQMAT”. M.C. acknowledges financial support from PNRR MUR project PE0000023-NQSTI. Author contributions G.S. fabricated the oxide heterostructures and the transport devices used in this study. G.S. and K.D. performed the measurements and analyzed the data with the support of S.G.. M.T.M., M.C., and C.O. developed the theoretical model. A.D.C. supervised the experimental work, and C.O. supervised the theoretical work. All authors contributed to writing the manuscript. Competing interests The authors declare no competing financial interests. Data and materials availability The data that support the findings of this study are available from Zenodo  (?) and from the corresponding authors upon reasonable request.
@@
Hall bar devices were fabricated by a combination of optical lithography and Ar ion milling (Fig. S1). The milling time was calibrated to stop the etching at the LaAlO3/SrTiO3 interface. Ti/Au contact pads were deposited by electron beam evaporation. The Hall bars were patterned along the $[\bar{1}10]$ and $[\bar{1}\bar{1}2]$ crystallographic directions of the substrate. The Hall bars have a width $W=20$ µm and a distance between the longitudinal probes $L=60$ µm.
@@
Devices along the orthogonal $[\bar{1}10]$ and $[\bar{1}\bar{1}2]$ directions were contacted with Al wire bonds. Electrically- and thermally-conductive Ag paint was used to anchor the samples to the chip carrier and apply a back-gate voltage. Four-probe magnetotransport measurements were performed in a liquid He4 Oxford Teslatron cryostat with base temperature of 1.7 K equipped with a superconducting magnet capable of generating magnetic fields up to 12 T. Nonlinear transport measurements were performed with the magnetic field applied in the sample plane along one of the two crystalline directions. Measurements of the ordinary Hall effect were performed with the magnetic field applied in the direction normal to the sample surface. The magnetic field was typically swept at a rate of 0.15 T/min. Commercial lock-in amplifiers (Zurich Instruments MFLI and Stanford Research 830) were used to apply an alternate current $I=50$ µA at a frequency $f=\frac{\omega}{2\pi}=17.7$ Hz and simultaneously detect the longitudinal and transverse harmonic voltages in both Hall bars. The amplitude of the electric current was controlled by setting the lock-in output voltage at a few Volts and by connecting a 100 kOhm resistor in series to the Hall bars. The actual current through the device was measured by terminating the circuit on the 50-Ohm current input of the lock-in amplifier. The DC back gate voltage was applied via a Keithley 2400 Sourcemeter. Prior to the measurements, an initial gate-forming procedure was followed to avoid hysteretic variations of the device resistance.
@@
The first harmonic longitudinal resistance has only a symmetric contribution, therefore $R^{\omega,\,\textrm{asym}}_{ii}\equiv R^{\omega}_{ii}$, with $i=x,\,y$. The first harmonic transverse resistance includes both symmetric and antisymmetric contributions. Since the symmetric part originates only from the contact misalignment and the anisotropic magnetoresistance, we neglect it and use the simplified notation $R^{\omega,\,\textrm{asym}}_{ji}=R^{\omega}_{ji}$. The second harmonic longitudinal resistance also includes symmetric and antisymmetric components. Since the nonlinear electronic transport driven by the quantum metric requires broken time reversal, only the antisymmetric part is considered in the main text. The symmetric component is discussed in the supplementary material.
The first and second harmonic resistances were calculated from the measured voltages as $R^{\omega}_{ji}=\frac{V^{\omega}_{ji}}{I_{i}^{\omega}}$ and $R^{2\omega}_{jii}=\frac{V^{2\omega}_{jii}}{I_{i}^{\omega}}$, respectively, with $i,j=x,\,y$. $i$ is the direction of the injected current, and $j$ is the direction of the measured voltage. Longitudinal (transverse) signals correspond to $i=j$ ($i\neq j$). The sheet conductivity was calculated from the first harmonic longitudinal resistance as $\sigma_{ii}=\frac{L}{WR^{\omega}_{ii}}$. The nonlinear conductivity $\sigma^{2\omega}_{iii}$ was defined as $J^{2\omega}_{i}=\sigma^{2\omega}_{iii}(E_{i}^{\omega})^{2}=\sigma_{ii}E_{i}^{2\omega}$, where $J$ and $E$ are the current density and electric field, respectively. The longitudinal nonlinear conductivity was therefore obtained from the relation $\sigma^{2\omega}_{iii}=\frac{R^{2\omega}_{iii}}{(R^{\omega}_{ii})^{3}I_{i}^{\omega}}\frac{L^{2}}{W}$.
@@ 1.1 The quantum geometric tensor of a Bloch wave
The quantum geometric tensor (QGT) characterizing the cell-periodic part of a Bloch wave $|u_{n}({\mathbf{k}})\rangle$ that satisfies the Schrödinger equation ${\mathcal{H}}({\mathbf{k}})|u_{n}({\mathbf{k}})\rangle=\varepsilon_{n}({\mathbf{k}})|u_{n}({\mathbf{k}})\rangle$, with ${\mathcal{H}}({\mathbf{k}})$ the Bloch Hamiltonian and $n$ a band index, can be introduced by recalling that treating the $d$-dimensional crystalline momentum ${\mathbf{k}}$ as a parameter, the distance between two quantum states in parametric space is defined as (15,54);
@@ 1.2 Nonlinear magnetotransport in a Rashba two-dimensional electron system
$\Omega_{\pm}$ $=$ $\mp\frac{1}{2}\hat{\mathbf{h}}\cdot(\partial_{k_{x}}\hat{\mathbf{h}}\times\partial_{k_{y}}\hat{\mathbf{h}})\;,$ (17)
where (±)plus-or-minus($\pm$)( ± ) are the band indices, and $a$ and $b$ are Cartesian coordinates ($x,y$). For the simple Rashba Hamiltonian in Eq. 14 the Berry curvature is zero, while for the quantum metric we have
@@
[…]
@@
[…]
@@ 3 Ordinary Hall effect
Figure S5 shows the ordinary Hall effect measured while sweeping the magnetic field in the direction normal to the sample surface. The nonlinearity of the Hall resistance with respect to the magnetic field indicates the presence of two bands at high gate voltage and/or low temperature. We use the two-band model (Methods) to determine the carrier density and mobility of each band as shown in Fig. 3 of the main text. We find that the carrier density and mobility along the $[\bar{1}10]$ and $[\bar{1}\bar{1}2]$ crystallographic directions differ by less than 10% and show the same dependence on temperature and gate voltage. This suggests a weak dependence of the transport properties on the crystallographic direction in the experimentally accessible range of chemical potential. This observation matches with the phenomenology of the measurements with a planar magnetic field.
@@ 4 Linear magnetoresistance
When the field is oriented out of the plane, the magnetoresistance is determined by the combination of the ordinary Lorentz magnetoresistance and weak (anti)localization effects (55,56). The analysis of this magnetoresistance allows us to extract the strength of the Rashba coupling, as discussed in Supplementary Section 5. The details of the analysis can be found in Ref. (6).
@@ 10 Exclusion of spurious effects
Figure 14: Fig. S10. Nonlinear electronic transport with $\mathbf{B}\parallel\mathbf{y}$. (A-B) Second harmonic longitudinal $B$-antisymmetric resistance measured as a function of the magnetic field $B$ and gate voltage $V_{\textrm{g}}$ for two relative orientations of $\mathbf{B}\parallel\mathbf{y}$ (black arrow) and the electric current $I$ (red arrow) at a temperature $T=3$ K. The electric current is applied along the $\bar{1}10$ (x) and $\bar{1}\bar{1}2$ (y) crystallographic directions in A and B, respectively. (C) Nonlinear longitudinal conductivity calculated from A as a function of the linear conductivity. (D) Dependence of the nonlinear conductivity on the square of the mobility of the second Rashba pair. The lines are linear fits to the data. (E) Magnetic field dependence of the quantum metric and bilinear magnetoresistance contributions to the total nonlinear conductivity obtained from the fits in D. The shaded areas correspond to the error of the fits.
Figure 15: Fig. S11. Nonlinear transport at variable temperature. (A-C) First harmonic transverse and second harmonic longitudinal resistances as a function of the in-plane magnetic field $B$ and temperature at a gate voltage $V_{\textrm{g}}=20$ V. The black and red arrows indicate the direction of $B$ and the electric current $I$, respectively. The current is applied along the $\bar{1}10$ $(x)$ and $\bar{1}\bar{1}2$ $(y)$ crystallographic directions in B and A, C, respectively. (D) Nonlinear longitudinal conductivity calculated from C at different magnetic fields as a function of the zero-field longitudinal linear conductivity. (E) Two-band electron densities $n$ and mobilities $\mu$ at variable temperature and a gate voltage $V_{\textrm{g}}=20$ V. (F) Dependence of the nonlinear conductivity on the square of the electron mobility of the second Rashba pair.

### v3

@@
Finally, we have performed a similar scaling analysis of the anomalous planar Hall resistance to account for the presence of disorder-mediated effects beside the intrinsic Berry curvature-induced contribution. The linear dependence of the anomalous planar Hall conductivity on the mobility $\mu_{2}$ (Fig. 4A) demonstrates the existence of a skew scattering contribution $\sigma_{\textrm{sk}}$ that is absent in a rotational symmetric Rashba two-dimensional electron system  (?) but is activated by the nonvanishing Berry curvature triple owing to hexagonal warping effects  (?). In addition, the $\tau^{0}$ term due to the combined action of side-jump and intrinsic Berry curvature varies nonmonotonically as the magnetic field is increased. This behavior is correctly captured by a theory calculation of the Berry curvature-mediated anomalous planar Hall response (Fig. 4D), which indicates that the intrinsic contribution dominates the $\tau$-independent anomalous planar Hall conductivity. Therefore, the intrinsic contributions to the anomalous planar Hall effect and nonlinear in-plane magnetoresistance provide us with direct access to the Berry curvature and the quantum metric of the 111-oriented LaAlO3/SrTiO3 interface and allow us to probe electronic dynamics associated with its full quantum geometric tensor. This unique property marks the difference of 111-oriented LaAlO3/SrTiO3 with respect to $\mathcal{P}\mathcal{T}$-symmetric materials like topological antiferromagnets  (?, ?), where the quantum geometric tensor is purely real and accessible only via the nonlinear response.
Discussion and outlook
@@
The analysis of the conductivity dependence on the scattering time allowed us to identify the intrinsic contribution of the quantum metric and Berry curvature to the electronic magnetotransport. Whereas the Berry curvature is enabled by the trigonal symmetry of the 111-oriented interface, the quantum metric is rotationally invariant and is not contingent on specific symmetry requirements beside the broken inversion symmetry. This property makes quantum metric effects potentially observable in a vast class of nonmagnetic materials. Moreover, the common origin of the QMMR and BMER suggests that all platforms in which the BMER was previously reported also host the quantum metric. The QMMR is thus not limited to oxide heterostructures  (?, ?) but is expected for all materials with spin-locked electronic bands such as semiconductors  (?, ?, ?, ?), surface states of heavy metals  (?), magnetic and nonmagnetic interfaces  (?, ?, ?), and even policrystalline samples. Our findings, therefore, do not only demonstrate the existence of a quantum metric associated with the spin-momentum locking but suggest new directions for investigating and functionalizing the quantum geometry. Besides magnetotransport, the quantum metric of the spin-momentum locking is, in principle, accessible to THz photocurrents and spectroscopic tools in time-reversal symmetric conditions   (?, ?). From a technological perspective, integrating the spin-momentum locking with magnetic and ferroelectric thin films could lead to the on-demand, reversible, and nonvolatile control of the physical effects emerging from the quantum metric. Moreover, combining spintronics with the quantum metric may reveal unforeseen disorder-free mechanisms to interconvert charge and spin-orbital currents  (?, ?, ?). This study, therefore, may enable a range of possibilities to harness quantum metric effects in condensed matter and may ultimately inspire technologies based on the geometry of the electronic wavefunctions.
@@ ACKNOWLEDGMENTS
We thank Jean-Marc Triscone, Marc Gabay, Ajit Srivastava, Edouard Lesne, and Graham Kimbell for fruitful discussions. We thank Marco Lopes and Jean-Luc Lorenzoni for technical support. We thank Alberto Morpurgo and Ignacio Gutiérrez for contributing to the sample fabrication. Funding: This work was supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract no. MB22.00071, by the Gordon and Betty Moore Foundation (grant no. 332 GBMF10451 to A.D.C.), by the European Research Council (ERC), and by the Dutch Research Council (NWO) as part of the VIDI (project 016.Vidi.189.061 to A.D.C.). G.S. acknowledges support from the Swiss National Science Foundation (grant no. PZ00P2_223542). A.D.C, M.T.M and M.C. acknowledge support from the EU Horizon 2020 research and innovation program under Grant Agreement No. 964398 (SUPERGATE). M.T.M. and C.O. acknowledge support from the MAECI project ”ULTRAQMAT” and the PNRR MUR project PE0000023-NQSTI (TOPQIN). M.C. acknowledges financial support from PNRR MUR project PE0000023-NQSTI. Author contributions: G.S. fabricated the oxide heterostructures and the transport devices used in this study. G.S. and K.D. performed the measurements and analyzed the data with the support of S.G.. M.T.M., M.C., and C.O. developed the theoretical model. A.D.C. supervised the experimental work, and C.O. supervised the theoretical work. All authors contributed to writing the manuscript. Competing interests: The authors declare no competing financial interests. Data and materials availability: The data and code that support the findings of this study are available from Zenodo  (?).
@@
Hall bar devices were fabricated by a combination of optical lithography and Ar ion milling (Fig. S1). The milling time was calibrated to stop the etching at the LaAlO3/SrTiO3 interface. Ti/Au contact pads were deposited by electron beam evaporation. The Hall bars were patterned along the $\bar{1}10]$ and $\bar{1}\bar{1}2]$ crystallographic directions of the substrate. The Hall bars have a width $W=20$ µm and a distance between the longitudinal probes $L=60$ µm.
@@
Devices along the orthogonal $\bar{1}10]$ and $\bar{1}\bar{1}2]$ directions were contacted with Al wire bonds. Electrically- and thermally-conductive Ag paint was used to anchor the samples to the chip carrier and apply a back-gate voltage. Four-probe magnetotransport measurements were performed in a liquid He4 Oxford Teslatron cryostat with base temperature of 1.7 K equipped with a superconducting magnet capable of generating magnetic fields up to 12 T. Nonlinear transport measurements were performed with the magnetic field applied in the sample plane along one of the two crystalline directions. Measurements of the ordinary Hall effect were performed with the magnetic field applied in the direction normal to the sample surface. The magnetic field was typically swept at a rate of 0.15 T/min. Commercial lock-in amplifiers (Zurich Instruments MFLI and Stanford Research 830) were used to apply an alternate current $I=50$ µA at a frequency $f=\frac{\omega}{2\pi}=17.7$ Hz and simultaneously detect the longitudinal and transverse harmonic voltages in both Hall bars. The amplitude of the electric current was controlled by setting the lock-in output voltage at a few Volts and by connecting a 100 kOhm resistor in series to the Hall bars. The actual current through the device was measured by terminating the circuit on the 50-Ohm current input of the lock-in amplifier. The DC back gate voltage was applied via a Keithley 2400 Sourcemeter. Prior to the measurements, an initial gate-forming procedure was followed to avoid hysteretic variations of the device resistance.
@@
The first harmonic longitudinal resistance has only a symmetric contribution, therefore $R^{\omega,\,\textrm{asym}}_{ii}\equiv R^{\omega}_{ii}$, with i=x,yi=x,\,y. The first harmonic transverse resistance includes both symmetric and antisymmetric contributions. Since the symmetric part originates only from the contact misalignment and the anisotropic magnetoresistance, we neglect it and use the simplified notation $R^{\omega,\,\textrm{asym}}_{ji}=R^{\omega}_{ji}$. The second harmonic longitudinal resistance also includes symmetric and antisymmetric components. Since the nonlinear electronic transport driven by the quantum metric requires broken time reversal, only the antisymmetric part is considered in the main text. The symmetric component is discussed in the supplementary material.
The first and second harmonic resistances were calculated from the measured voltages as $R^{\omega}_{ji}=\frac{V^{\omega}_{ji}}{I_{i}^{\omega}}$ and $R^{2\omega}_{jii}=\frac{V^{2\omega}_{jii}}{I_{i}^{\omega}}$, respectively, with i,j=x,yi,j=x,\,y. $i$ is the direction of the injected current, and $j$ is the direction of the measured voltage. Longitudinal (transverse) signals correspond to $i=j$ ($i\neq j$). The sheet conductivity was calculated from the first harmonic longitudinal resistance as $\sigma_{ii}=\frac{L}{WR^{\omega}_{ii}}$. The nonlinear conductivity $\sigma^{2\omega}_{iii}$ was defined as $J^{2\omega}_{i}=\sigma^{2\omega}_{iii}(E_{i}^{\omega})^{2}=\sigma_{ii}E_{i}^{2\omega}$, where $J$ and $E$ are the current density and electric field, respectively. The longitudinal nonlinear conductivity was therefore obtained from the relation $\sigma^{2\omega}_{iii}=\frac{R^{2\omega}_{iii}}{(R^{\omega}_{ii})^{3}I_{i}^{\omega}}\frac{L^{2}}{W}$.
@@ 1.1 The quantum geometric tensor of a Bloch wave
The quantum geometric tensor (QGT) characterizing the cell-periodic part of a Bloch wave $|u_{n}({\mathbf{k}})\rangle$ that satisfies the Schrödinger equation ${\mathcal{H}}({\mathbf{k}})|u_{n}({\mathbf{k}})\rangle=\varepsilon_{n}({\mathbf{k}})|u_{n}({\mathbf{k}})\rangle$, with ${\mathcal{H}}({\mathbf{k}})$ the Bloch Hamiltonian and $n$ a band index, can be introduced by recalling that treating the $d$-dimensional crystalline momentum $\mathbf{k}$} as a parameter, the distance between two quantum states in parametric space is defined as (15,56);
@@ 1.2 Nonlinear magnetotransport in a Rashba two-dimensional electron system
[…]
where $(\pm)$ are the band indices, and $a$ and $b$ are Cartesian coordinates ($x,y$). For the simple Rashba Hamiltonian in Eq. 14 the Berry curvature is zero, while for the quantum metric we have
@@
$\sigma_{xxx}^{(\rm{qm})}(B<B_{\textrm{c}})$ $=$ $\frac{e^{3}}{\hbar}\frac{1}{(2\pi)^{2}}\int_{0}^{2\pi}d\phi\int_{p_{+}^{(+)}(\phi)}^{p_{+}^{(-)}(\phi)}d\,p\frac{15\cos\phi\sin^{2}\phi}{16|\alpha_{R}|p^{3}}$ (27)
$=$ $\frac{e^{3}}{\hbar}\frac{15}{16|\alpha|}\frac{1}{(2\pi)^{2}}\int_{0}^{2\pi}d\phi\cos\phi\sin^{2}\phi\frac{1}{2}\left[\frac{1}{(p_{+}^{(+)}(\phi))^{2}}-\frac{1}{(p_{+}^{(-)}(\phi))^{2}}\right]$
$=$ $\frac{e^{3}}{\hbar}\frac{15}{16|\alpha|}\frac{1}{(2\pi)^{2}}\int_{0}^{2\pi}d\phi\cos\phi\sin^{2}\phi\left[g_{1}(\phi)+g_{2}(\phi)\right]\;,$
@@
$\sigma_{xxx}^{(\rm{bm})}$ $=$ $\tau^{2}\frac{e^{3}}{\hbar^{3}}\frac{1}{(2\pi)^{2}}\int d\phi d\,p\frac{3|\alpha_{R}|\cos\phi\sin^{2}\phi}{p}(f_{+}(\mathbf{p})-f_{-}(\mathbf{p}))$ (30)
$\sigma_{xxx}^{(\rm{bm})}(B<B_{\textrm{c}})$ $=$ $\tau^{2}\frac{e^{3}}{\hbar^{3}}\frac{3|\alpha_{R}|}{(2\pi)^{2}}\int_{0}^{2\pi}d\phi\;\cos\phi\sin^{2}\phi\;\ln{\left(\frac{p_{+}^{(-)}(\phi)}{p_{+}^{(+)}(\phi)}\right)}=0$ (31)
@@ 3 Ordinary Hall effect
Figure S5 shows the ordinary Hall effect measured while sweeping the magnetic field in the direction normal to the sample surface. The nonlinearity of the Hall resistance with respect to the magnetic field indicates the presence of two bands at high gate voltage and/or low temperature. We use the two-band model (Methods) to determine the carrier density and mobility of each band as shown in Fig. 3 of the main text. We find that the carrier density and mobility along the $\bar{1}10]$ and $\bar{1}\bar{1}2]$ crystallographic directions differ by less than 10% and show the same dependence on temperature and gate voltage. This suggests a weak dependence of the transport properties on the crystallographic direction in the experimentally accessible range of chemical potential. This observation matches with the phenomenology of the measurements with a planar magnetic field.
@@ 4 Linear magnetoresistance
When the field is oriented out of the plane, the magnetoresistance is determined by the combination of the ordinary Lorentz magnetoresistance and weak (anti)localization effects (57,58). The analysis of this magnetoresistance allows us to extract the strength of the Rashba coupling, as discussed in Supplementary Section 5. The details of the analysis can be found in Ref. (6).
@@ 10 Exclusion of spurious effects
Figure 14: Fig. S10. Nonlinear electronic transport with $\mathbf{B}\parallel\mathbf{y}$. (A-B) Second harmonic longitudinal $B$-antisymmetric resistance measured as a function of the magnetic field $B$ and gate voltage $V_{\textrm{g}}$ for two relative orientations of $\mathbf{B}\parallel\mathbf{y}$ (black arrow) and the electric current II (red arrow) at a temperature $T=3$ K. The electric current is applied along the $\bar{1}10$ (x) and $\bar{1}\bar{1}2$ (y) crystallographic directions in A and B, respectively. (C) Nonlinear longitudinal conductivity calculated from A as a function of the linear conductivity. (D) Dependence of the nonlinear conductivity on the square of the mobility of the second Rashba pair. The lines are linear fits to the data. (E) Magnetic field dependence of the quantum metric and bilinear magnetoresistance contributions to the total nonlinear conductivity obtained from the fits in D. The shaded areas correspond to the error of the fits.
Figure 15: Fig. S11. Nonlinear transport at variable temperature. (A-C) First harmonic transverse and second harmonic longitudinal resistances as a function of the in-plane magnetic field $B$ and temperature at a gate voltage $V_{\textrm{g}}=20$ V. The black and red arrows indicate the direction of $B$ and the electric current II, respectively. The current is applied along the $\bar{1}10$ $(x)$ and $\bar{1}\bar{1}2$ $(y)$ crystallographic directions in B and A, C, respectively. (D) Nonlinear longitudinal conductivity calculated from C at different magnetic fields as a function of the zero-field longitudinal linear conductivity. (E) Two-band electron densities $n$ and mobilities $\mu$ at variable temperature and a gate voltage $V_{\textrm{g}}=20$ V. (F) Dependence of the nonlinear conductivity on the square of the electron mobility of the second Rashba pair.